the first three terms of a geometric sequence are: T1:T2:T3. if T2=T1+4 and T3=T2+9, determine the value of T1:T2:T3
Answer by fractalier(6550)

If that is so, then the ratio of T2 to T1 = the ratio of T3 to T2, or
T2/T1 = T3/T2
T2/(T2-4) = (T2+9)/T2
Let us call T2, x, for simplicity...now cross-multiply and get
x^2 = (x-4)(x+9)
x^2 = x^2 + 5x - 36
0 = 5x - 36
36 = 5x
x = T2 = 7.2
T1 = 3.2
T3 = 16.2
so that
T1:T2:T3 = 3.2:7.2:16.2 = 16:36:81